OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, -1) and has a slope of 5. y - 1 = 5(x - 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y - 2 = 5(x - 1)

OpenStudy (luigi0210):

Point-slope formula: y-y1=m(x-x1) --------------------------- m=slope y1=-1 x1=-2 ---------------------------

OpenStudy (anonymous):

so the answer is y - 1 = 5(x - 2)?

OpenStudy (luigi0210):

Ah, watch your negatives.

OpenStudy (luigi0210):

\[y-(-1)=5(x-(-2))\]

OpenStudy (anonymous):

so what will the answer be then?

OpenStudy (luigi0210):

>.>

OpenStudy (anonymous):

the one i said? or was i wrong u_u

OpenStudy (luigi0210):

what's -1*-1?

OpenStudy (anonymous):

its 1

OpenStudy (luigi0210):

so: y-(-1) =y+1

OpenStudy (anonymous):

y + 1 = 5(x + 2) ? that'll be the answer

OpenStudy (luigi0210):

There you go :P